RHIC Startup:
Increasing bunches through 12x12, 28x28 to 37x37 giving ZDC rates of ~15kHz at both STAR and PHENIX and beam currents of 7E10 deuterons and 6E8 gold. Single beam (Blue) running with 60 and 110 bunches. Today go to colliding 56x56 first. Then work on fixing the beam radius. By midnight tonight, we could have collisions of 56x56. The experiments can start timing in their detectors.
Continued vacuum work in AGS, this time at L07 where replacing a flange on a blank-off valve seems to have stopped the leak for the moment. Ion pump at L07 tripped off. Bad vacuum was responsible for RF tripping off.
Feedback and replay working for Blue. Ramp efficiency at 90% for bleu with ARTUS off and feedback on, somewhat less for Yellow. No feedback yet for Yellow. There appears to be an overlap with a 60 Hz resonance and the Yellow tune. The Yellow tune needs to be changed in the Ramp, although Larry offered to put it in by hand so that work on Yellow could go forward. Steve Tepekian, Jim Cameron, Ron Schroeder will work on getting feedback working for Yellow by getting BBQ working, improving the tunes in Yellow to avoid 60Hz overlap.
Angelika needs to work on the collimators.
Cavity conditioning, 48 hour estimate, was overly conservative. We are in a position to start rebucketing tomorrow.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email:sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |